OpenStudy (anonymous):

Algebra Help: 1. Simplify Which is equal to 4x^2 + x - 7 / x^2 - 7 + 4x - 3x^2 / x^2 - 7 A. 8x^2 - 2x - 7 / x^2 - 7 <-- it's either A B. 7x^2 - 3x - 7 / x^2 - 7 C. x^2 + 5x - 7 / x^2 - 7 <-- or it's either C D. x^2 - x - 7 / x^2 - 7

OpenStudy (tkhunny):

You will have to remember to use parentheses when necessary. This is what you have written: 4x^2 + x - 7 / x^2 - 7 + 4x - 3x^2 / x^2 - 7 = \(4x^{2} + x - \dfrac{7}{x^2} - 7 + 4x - \dfrac{3x^{2}}{x^2} - 7\) I'm going to guess this is not what you intended.